OpenStudy (anonymous):

Graph the line with the slope 1/3 passing through the point( -2,5)

OpenStudy (anonymous):

y=mx+b?

pooja195 (pooja195):

Not thats slope intercept :3 \[\huge y-y1=m(x-x1)\]

pooja195 (pooja195):

\[\huge(x1,y1)\] \[\huge~m=slope\]

OpenStudy (anonymous):

so m would be 1/3 and x1 y1 would be -2,5?

pooja195 (pooja195):

yep and now we plug it in :) \[\huge~y-5=\frac{ 1 }{ 3 }(x+2)\] two negatives make a postitve thats where the 2 came from :) can u solve this now?

OpenStudy (anonymous):

would the answer be m=1/3 and b= 17-3?

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

\[\huge y=\frac{ 1 }{ 3 }x+\frac{ 17 }{3 }\]

pooja195 (pooja195):

Now u just graph that

OpenStudy (anonymous):

uhmm how exactly would I graph that

pooja195 (pooja195):

use this https://www.desmos.com/calculator

OpenStudy (anonymous):

Thank you so much :)

pooja195 (pooja195):

:)